OpenStudy (anonymous):

solve. 7^(x – 5) = 7^(4x + 7). a. x = –4 b. x = -12/5 c. x = 2/3 d. x = 4

OpenStudy (anonymous):

a.

sam (.sam.):

7^(x-5)=7^(4x+7) (x-5)=(4x+7) x-5=(4x+7) x-5=4x+7 -3x-5=7 -3x=12 x=-4

OpenStudy (anonymous):

I am giving @.Sam. a medal for the solution process.